OpenStudy (anonymous):

1.5 g of a gaseous compound of carbon and hydrogen occupies 1.12 dm^3 at STP. What is the mass of 1 mole of this compound and what is its likely formula? Thanks!! will give medal!

OpenStudy (preetha):

You have been given volume, and STP. Soyou can calculate moles of this compound. 1 mole of any gas occupies 22.4 mL at STP. Calculate the moles and then convert that to molar mass. Moles = g/molar mass. Then you will have to look at the molar mass and go some guessing.

OpenStudy (anonymous):

i got the No. of moles which is 0.05 mol and the molar mass which is 30 g/mol. Im stuck there. I don't understand what to do to find the formula. please help.

thomaster (thomaster):

gaseous compound of carbon and hydrogen so \(\to \sf\Large C_xH_y\) 1 mol gas occupies \(\sf 22.4\;dm^3\;(or\;Liter)\\\) So \(\sf\frac{1.12}{22.4}=0.05\;mol\\\) \(\sf\frac{1.5}{0.05}=30\;g/mol\) \(\sf \to(This~it~the~point~where~you~are~now)\gets\) \(\Large{\hspace{4pt}\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow~\downarrow}\\\) Now you have to guess which \(\sf C_xH_y\) gas has a molecular weight of 30 g/mol. Carbon = 12 g/mol (actually it's 12.011 but we can use rough numbers) Hydrogen = 1 g/mol First one you'd try is methane \(\sf \large{CH_4}\;\to \normalsize{12+(4*1)}=16 g/mol\) Next possible one is ethane (because the general form is \(\sf C_nH_{n+2}\) \(\sf \large{C_2H_6}\;\to \normalsize{(2*12)+(6*1)}=30 g/mol\) Now we know that the compound is ethane \(\Large\ddot\smile\) \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)

thomaster (thomaster):

I made a slight mistake, the general form of a hydrocarbon is not \(\large\sf C_nH_{\underline{n}+2}\) its \(\large\sf C_nH_{\underline{2n}+2}\)

OpenStudy (anonymous):

Thanks alot!!!!!!! it helped:)

thomaster (thomaster):

I'm glad it did \(\Large\ddot\smile\)